# BlockLLM: Multi-tenant Finer-grained Serving for Large Language Models

### v1

@@ 4.2 Model Partitioning
With the evaluated equivalence, BlockLLM incorporates an automated mechanism that breaks down LLMs into blocks. Each block then becomes an independent engine during serving.
@@
As for FF models, after BlockLLM evaluates equivalence between model components ( 4.1), we perform lazy partitioning. As shown in Figure 11, for a FF model adapted from the foundation model (step 1), BlockLLM sets a threshold $T$ and group all connected network components (e.g., transformer layer 1 from the fine-tuned model and the foundation) that have equivalence exceeding $T$ into a block (step 2). This action results in 4 blocks, with two blocks being considered equivalent. When a newly added PE model (e.g., LoRA) arrives (step 3), BlockLLM (1) preserve its PE adapters (e.g., AttLoRA1, AttLoRA2), (2) find all existing blocks that share parameters (e.g., FFN1, FFN2) with the new model and (3) attempts to partition the existing blocks to enable block reuse. LoRA here has a different Attention layer. Therefore, we further partition the foundation model into four blocks so that they share the FFN layers in separately-provision blocks. This action results in 9 blocks with two chain-of-blocks being equivalent.
@@ 4.3 Stitching Block
Now that BlockLLM has partitioned LLMs into blocks, we tackle the third question. When two blocks of different embedding sizes are considered equivalent, such as Blocks B and D illustrated in Figure 9, how to route requests originally intended for, say Block B to D, becomes a unique challenge.
@@
We use a Linear layer to serve as the stitching block. This stitching block is trained while maintaining the parameters of other blocks static. To generalize the stitching block, we encode the positional information of the stitching point as an extra dimension. The position value is the sum of the positions of the head block and the tail block in its original foundation model. During training, it is initially placed at a shallow stitchable layer and progressively moved to deeper ones if there is any. This approach follows the idea of reusing layers in [22, 41]. Table 3 shows the training costs and the performance of the stitched model compared with the large model.
@@ 5.1 KV Cache Coordination
Memory bandwidth-bound KV cache. Efficient stateful coordination of the KV cache is critical for auto-regressive LLMs in BlockLLM because memory bandwidth constraints on the KV cache emerge as a significant bottleneck, as evidenced by plenty of studies. Existing systems serve one batch of requests at a time and thus only consider the trade-off between recalculating the KV matrices and the cost of caching the KV matrices in the device memory. This trade-off reaches a point— determined by factors such as device type, model architecture, and request sequence length—where caching becomes more efficient than recalculation. However, as request sequences become longer, memory bandwidth constraints increasingly invariably become a performance-bounding factor when loading the KV cache [18].
@@
I/O and recalculation cost. As aforementioned in 3.2, BlockLLM’s design complicates the problem. The assumption that requests will consistently be processed by the same block instances throughout all iterations no longer holds. To exploit the KV cache, the I/O cost of transferring cache between block instances may be unavoidable. We thus examine the cost implications within BlockLLM’s design space, considering two potential scenarios when a request batch currently residing on device $d_{i}$ initiates a new iteration.
In the first scenario, the request revisits the block instance on device $d_{j}$ that retains its KV cache. The latency is thus influenced by P2P request transfer overhead and the loading cost of KV cache. The increased latency can be expressed as:
$T_{transfer}^{w/kv}=\frac{D^{\prime}_{req}}{B_{net}(d_{i},d_{j})}+\frac{D_{cache}}{B_{mem}(d_{j})},$
where $D^{\prime}_{req}$ is the size of the newly-generated token and $D_{cache}$ is the size of its KV cache. $B_{net}(d_{i},d_{j})$ is the network bandwidth between two devices and $B_{mem}(d_{j})$ is the device memory bandwidth of $d_{j}$.
In the second scenario, the request is dispatched to a different block instance on device $k$ without the memory of its KV cache. The increased latency encompasses P2P request transfer overhead and either the P2P transfer cost of the KV cache or the recalculation cost of the KV cache. The resulting increased latency is formulated as:
$\begin{split}T_{transfer}^{w/kv}&=\min(\frac{D^{\prime}_{req}}{B_{net}(d_{i},d_{k})}+\frac{D_{cache}}{B_{net}(d_{j},d_{k})}+\frac{D_{cache}}{B_{mem}(d_{k})},\\ &\frac{D_{req}}{B_{net}(d_{i},d_{k})}+\frac{D_{cache}}{B_{comp}(d_{k})}),\end{split}$
where $D_{req}$ encapsulates the entire request, including the original input and all subsequently generated tokens. $B_{comp}(d_{k})$ denotes the computation capacity (i.e. FLOPS) on device $k$. We consider two cases for KV cache handling: recalculating the KV cache or transferring it. The former entails the transfer of both the request and its associated KV cache to the target block instance and the cost of loading the KV cache during computation, while the latter involves the transfer cost of the complete sequence for recalculation of the KV matrices.
Prioritize KV cache owner. Comparing these two scenarios, it becomes evident that transferring the KV cache to a new block instance is consistently less efficient than returning to the original KV cache owner. As for recalculating KV on the new block instance, it is also less efficient because caching typically outperforms recalculations, and network bandwidth is generally not as high-capacity as memory bandwidth. We follow this insight and perform best-effort coordination. As long as the candidate block instances have the same status (i.e., queuing time), BlockLLM’s agent should give priority to dispatch the request to the block instance that holds the associated KV cache. In 5.3, we discuss in detail how BlockLLM’s agent picks the suitable block instance.
Ownership of KV cache. BlockLLM is designed to store a KV cache for multiple requests, making it essential to distinguish the ownership of each KV cache matrix. Each agent maintains a dictionary mapping requests to the addresses of their corresponding KV caches. When processing a batch of requests, an agent of BlockLLM looks up this dictionary using the hashed request ID to retrieve the memory address of the relevant KV matrices. Additionally, multiple devices can hold a KV cache for a single request. To avoid this redundancy, the most recent copy is retained and BlockLLM’s scheduler periodically removes the others. When BlockLLM’s scheduler receives a completed request, it lets the agents possessing the KV cache of that request delete the associated memory.
@@ 5.2 Speculative Execution
Request transfer overhead hurts the overall throughput in two ways. First, it directly and inevitably prolongs inference latency due to its inherent blocking characteristic: each subsequent block’s operation is contingent upon receiving output from the preceding block, creating a strong dependency. Second, it exhausts network resources; each redirection of a request to a different server incurs network involvement. Moreover, as BlockLLM accommodates larger batch sizes than traditional serving systems, the size of data transferred increases, consequently consuming more resources. In BlockLLM, we propose to tackle the first with speculative execution and the second with block scheduling and placement.
@@
Block surrogates. We leverage insights from existing work and empirical studies, coupled with our block zoo, to construct high-fidelity surrogates for the blocks. The surrogates are used for prediction. First, we consider pruned models. One typical pruning technique is sparse models. While sparsity typically results in a reduction of computation FLOPs, it does not always correspond to a commensurate decrease in computation time. Hence, our focus is on pruning techniques that meaningfully cut down computation time. We implement a strategy inspired by existing literature [23] to construct surrogates for our blocks. In particular, we perform structured pruning by selectively removing sub-structure that generates less critical impact towards the model output and then add fine-tuned LoRA for performance recovery after the pruning process. Table 4 shows the output similarity and computation time of four blocks and their surrogates. Second, we exploit equivalent blocks with different computation costs in BlockLLM’s block zoo but in different computation costs in BlockLLM’s zoo. For instance, a block partitioned from a 7B LLaMA model can serve as an effective surrogate for a block from a 30B LLaMA model, which is adopted by [26].
@@
Bottleneck-pinpointing speculation. Naively application of speculative execution to blocks is not a guarantee of latency improvement due to the necessity of verifying predicted results. In Figure 13, consider an inference pipeline composed of three blocks where speculative execution is applied to each. Ideally, if all speculations are destined to be correct, the inference latency is expected to be reduced from T6 to T2 ( 1). However, consecutive prediction verification can diminish these gains, especially when erroneous predictions happen at the latter part of the chain. For instance, if a prediction error arises at Block C (last block in the chain) ( 4), the latency reduction would only be 0, compared to a reduction of T6-T3 if the error occurred earlier at Block A 2. Therefore, the optimal inference ( 5) one could achieve is T5 instead of T2. In the worst case, where all three speculations are incorrect, there would be no latency reduction at all. Practically, due to the concurrent execution of blocks and their surrogates, this could even result in additional slowdowns.
To avoid such outcomes, we follow two rules when implementing speculation. First, we avoid enabling speculation on consecutive blocks to minimize accumulated errors and wasted resources. Second, speculation is not applied at the last block in the chain because the final output, once relayed back to BlockLLM’s scheduler, is not correctable. Thus, during serving, we selectively apply speculation to the top-$k$ bottleneck blocks—those characterized by their computation intensity or frequent use.
@@ 5.3 Request and Block Scheduling
We then explain BlockLLM’s scheduling for requests and blocks.
@@
BlockLLM’s request scheduling strategy is to prioritize the KV cache owner instance as long as its device memory can hold the request data; otherwise estimate the latency of each candidate instance and schedule the request onto the instance with the lowest latency increase. When the KV cache owner is not available, the components of the latency estimation of a candidate block instance residing on device $c$ are as follows:
$\begin{split}Latency_{d_{c}}&=T_{queue}+T_{compute}+T_{transfer}+T_{load},\\ T_{queue}&=\sum_{i=1}^{n}{Comp(req_{i})},\\ T_{compute}&=Comp(req),\\ T_{transfer}&=\begin{cases}T_{transfer}^{w/kv}&\text{agents dispatch,}\\ \dfrac{D_{req}}{B_{net}(s,d_{c})}&\text{scheduler dispatches,}\end{cases}\\ T_{load}&=\begin{cases}0&\text{idle device,}\\ \dfrac{D_{b^{\prime}}}{B_{mem}(d_{c})}+\dfrac{D_{b}}{B_{net}(d_{c})}&\text{loading overhead.}\\ \end{cases}\end{split}$
We incorporate four key factors: queuing time, computation time, transfer time, and the overhead associated with block switching. Queuing time accounts for the duration required to process all $n$ queuing request batches. Transfer time is subject to who initiates the dispatch: if done by the scheduler $s$, the request batch is at the start of the chain and no KV cache is involved. Otherwise, agent dispatching includes the costs of transferring the request and its KV cache (discussed in 5.1). The overhead from switching block depends on the current status of the device; if idle, the block’s loading can be overlapped with other operations, thus zero overhead. Otherwise, this loading overhead includes the time to move out the existing block $b^{\prime}$ and load the new block $b$.
@@
In BlockLLM, the locality is quantified by monitoring historical traffic, specifically by a counter recording the frequency of being directly inter-dependent between two blocks. Block pairs with a higher locality value are placed within the same server. Furthermore, BlockLLM’s scheduler dynamically adapts to evolving traffic patterns; should the observed locality change, block instances are migrated as necessary to align with the new pattern.
6 Implementation
We have implemented a prototype of BlockLLM. It is compatible with HuggingFace models. We use NCCL as the communication backend to transfer data among servers.
Profiling. To support the online serving system, BlockLLM conducts profiling on blocks. First, we profile the computation time of each block. It includes the computation time measured under different batch sizes. This also applies to the surrogates of each block and the performance when the block and its surrogates are multiplexing. Second, we obtain the communication time from one device to another device in different transfer sizes using NCCL primitives. Third, we measure the overhead of loading the block engine from disk to host memory and to device memory and vice versa.
Batching. The computation efficiency improves with a larger batch size. However, forcing the blocks to adopt a fixed large batch size would complicate the problem because it involves reorganizing the requests from different batches/applications. Therefore, we loosely encourage batching within each block instance. Upon receiving a new batch of requests, BlockLLM’s agent puts it into the queue if there are any and attempts to pack it with its direct neighbors as long as the combined batch does not exceed the upper limits of the batch size of the block. If there is no queuing request, BlockLLM’s agent sends the request batch directly to the instance for processing. Additionally, in case one request in the batch reaches EOS at the last chain of block, it is removed from the batch and relayed to the scheduler for response.
Request dispatching. BlockLLM’s agents adopt a FIFO+priority queue, where priority is assigned to requests that have left KV cache memory and may revisit in the future. Each block instance maintains a countdown clock associated with every auto-regressive request based on the time estimation of one iteration, anticipating its potential return. As long as the countdown remains active, BlockLLM’s agent prioritizes the processing of returning requests.
The scheduler and agents dispatch the request in slightly different ways. BlockLLM’s agent needs to (1) determine the candidate block for each request and pack them in different buckets and (2) obtain the placement of the block candidates. BlockLLM’s agent initiates a broadcasting request, and those available agents hosting the candidates would respond. BlockLLM’s scheduler, on the other hand, keeps a live record of the placement of every block instance in the cluster, thus easing the dispatching process.

### v2

@@ 4.3 Model Partitioning
With the evaluated equivalence and accuracy, BlockLLM incorporates an automated mechanism that breaks down LLMs into blocks. Each block then becomes an independent engine during serving.
@@
As for FF models, after BlockLLM evaluates equivalence between model components (4.1), we perform lazy partitioning. As shown in Figure 12, for an FF model adapted from the foundation model ( 1), BlockLLM sets a threshold $T$ and group all connected network components (e.g., transformer layer 1 from the fine-tuned model and the foundation) that have equivalence exceeding $T$ into a block ( 2). This action results in 4 blocks, with two blocks being considered equivalent. When a newly added PE model (e.g., LoRA) arrives ( 3), BlockLLM (1) preserve its PE adapters (e.g., AttLoRA1, AttLoRA2), (2) find all existing blocks that share parameters (e.g., FFN1, FFN2) with the new model and (3) attempts to partition the existing blocks to enable block reuse. LoRA here has a different Attention layer. Therefore, we further partition the foundation model into four blocks so that they share the FFN layers in separately-provision blocks ( 4). This action results in 9 blocks with two chain-of-blocks being equivalent.
@@ 4.4 Stitching Block
Figure 13: Accuracy & TPOT after applying a stitching layer to LLaMA (1.1B, 7B, 13B)
While requests can be directly routed between blocks with the same embedding sizes, this does not apply to blocks of different embedding sizes. When two blocks of different embedding sizes are considered equivalent, such as Blocks B and D illustrated in Figure 9, routing requests originally intended for Block B to D presents a unique challenge.
@@
We use a Linear layer to serve as the stitching block. This stitching block is trained while keeping the parameters of other blocks static. This approach is inspired by [38]. To generalize the stitching block, we encode the positional information of the stitching point as an extra dimension. The position value is the sum of the positions of the head block and the tail block in their original foundation model. During training, the stitching block is initially placed at a shallow layer and progressively moved to deeper ones. This approach follows the idea of reusing layers in [29, 50]. Table 3 shows the training costs and the performance of the stitched model compared with the large model. Figure 13 shows the accuracy and latency when applying stitching block between different decoding layers in LLaMA 1.1B, 7B, and 13B.
@@ 5.1 KV Cache Coordination
Memory bandwidth-bound KV cache. Efficient stateful coordination of the KV cache is crucial for auto-regressive LLMs in BlockLLM, as memory bandwidth constraints on the KV cache have been identified as a significant bottleneck in numerous studies. Existing systems process one batch of requests at a time, weighing the trade-off between recalculating the KV matrices and caching them in device memory. This trade-off reaches a point—determined by factors such as device type, model architecture, and request sequence length—where caching becomes more efficient than recalculation. However, as request sequences lengthen, memory bandwidth constraints become a performance-bounding factor when loading the KV cache [25].
@@
I/O and recalculation cost. As aforementioned in 3.2, BlockLLM’s design complicates the problem. The assumption that requests are consistently processed by the same block instances no longer holds, making I/O costs for transferring KV caches between instances unavoidable.
To migrate the KV cache from device $d_{i}$ to $d_{j}$, we optimize the process by overlapping KV cache recomputation with copying, thereby minimizing migration time. Given the fully known context, we employ chunked prefilling for efficient recomputation. For sequences to be migrated, denoted as $S={s_{i}}_{1}^{n}$, we begin by recomputing the KV cache from the start of sequence $s_{1}$ while simultaneously copying the cache starting from the end of sequence $s_{n}$. The process concludes when recomputation encounters a KV cache page that has already been copied, indicating the completion of migration.
This approach is chosen for two key reasons. First, each token’s KV cache depends on the KV caches of all preceding tokens. Recomputing the KV cache from the beginning of a sequence ensures the accuracy of the entire cache. In contrast, copying can take place independently, without relying on preceding caches. Second, ongoing requests on the target device may introduce latency during recomputation. By employing chunked prefill, we improve GPU utilization and mitigate the impact of KV cache recomputation on other tasks.
Proactive KV Cache Migration. As BlockLLM may redirect requests to blocks lacking the necessary KV caches, this can introduce additional migration latency. While this overhead cannot be completely eliminated, it can be mitigated by proactively migrating KV caches in advance, thereby removing it from the critical path.
To ensure that migration does not introduce latency, it is essential to predict whether the KV cache will be used before the migration completes. The feasibility of predicting KV cache usage has been demonstrated in [4]. We adopt the method proposed in [4] to estimate the interception time: $T_{INT}=t_{now}-t_{call}$, where $t_{now}$ is the current time updated for each iteration, and $t_{call}$ is the time when the last interception was initiated.
Memory Efficiency. Modern LLM inference serving systems support paged attention [25], a technique that partitions the KV cache into smaller pages. This approach eliminates the need to store the entire KV cache in contiguous memory and allows for the sharing of KV cache pages across multiple requests, thereby enhancing memory efficiency. However, dynamically routing requests to blocks that lack the required KV cache can result in the creation of new KV cache pages. Since each device maintains its own dedicated KV cache page table, generating the same KV cache page on a different device leads to the duplication of KV pages. This duplication, which otherwise would only require a single KV page with an incremented reference counter, undermines the advantages of memory sharing.
To prevent memory waste, we prioritize migrating pages referenced by fewer requests before those referenced by more. We denote all KV pages as $C=\{c_{1},c_{2},...c_{n}\}$, where each $c_{i}$ represents the underlying consecutive KV pages of request $s_{i}$, and $n$ is the total number of requests tracked by the system. We use $ref(c_{i})$ to calculate the total number of pages referenced by more than one request in sequence $s_{i}$. For KV cache pages $c_{i}\in C$, we have $ref(c_{i})\leq ref(c_{i+1})$. If $ref(c_{i})=ref(c_{i+1})$, then $resumeTime(c_{i})<=resumeTime(c_{i+1})$, where $resumeTime(c_{i})$ is the estimated time the pages $c_{i}$ will be reused by the intercepted request $s_{i}$.
Prioritize KV cache owner. Transferring KV caches to a new block instance can introduce latency, impacting request processing times. This latency is subject to network bandwidth and varying network conditions. Although recalculating the KV cache can sometimes be more efficient than direct copying, it may still disrupt ongoing requests on the target device. Therefore, we prioritize block instances that already possess the required KV cache before redirecting requests to a new device. This approach follows the principle of best-effort coordination. When candidate block instances have the same status (e.g., queuing time), BlockLLM ’s agent prioritizes dispatching the request to the block instance that holds the associated KV cache. In 5.3, we provide a detailed discussion on how BlockLLM’agent selects the appropriate block instance.
@@ 5.2 Speculative Execution
Request transfer overhead negatively impacts overall throughput in two significant ways. First, it directly and inevitably increases inference latency due to its inherent blocking nature: each subsequent block’s operation depends on receiving output from the preceding block, creating a strong dependency. Second, it depletes network resources, as each redirection of a request to a different server involves network transfer. In BlockLLM, we address the first issue with speculative execution and the second with block scheduling and placement.
@@
Block surrogates. We leverage insights from existing work and empirical studies, coupled with our block zoo, to construct high-fidelity surrogates for the blocks. The surrogates are used for prediction. First, we consider pruned models. One typical pruning technique is sparse models. While sparsity typically results in a reduction of computation FLOPs, it does not always correspond to a commensurate decrease in computation time. Hence, our focus is on pruning techniques that meaningfully cut down computation time. We implement a strategy inspired by existing literature [30] to construct surrogates for our blocks. In particular, we perform structured pruning by selectively removing sub-structure that generates less critical impact towards the model output and then add fine-tuned LoRA for performance recovery after the pruning process. Table 4 shows the output similarity and computation time of four blocks and their surrogates. Second, we exploit equivalent blocks with different computation costs. For instance, a block partitioned from a 7B LLaMA model can serve as an effective surrogate for a block from a 30B LLaMA model, which is adopted by [34].
@@
Bottleneck-pinpointing speculation. Naively applying speculative execution to blocks does not guarantee latency improvement, as it requires verifying the predicted results. In Figure 15, consider an inference pipeline composed of three blocks where speculative execution is applied to each. Ideally, if all speculations are destined to be correct, the inference latency is expected to be reduced from T6 to T2 ( 1). However, consecutive prediction verification can diminish these gains, especially when erroneous predictions happen at the latter part of the chain. For instance, if a prediction error arises at Block C (last block in the chain) ( 4), the latency reduction would only be 0, compared to a reduction of T6-T3 if the error occurred earlier at Block A ( 2). Therefore, the optimal inference ( 5) one could achieve is T5 instead of T2, because each speculative block needs to wait for the completion of the bigger verification block. In the worst case, where all three speculations are incorrect, there would be no latency reduction at all. Practically, due to the concurrent execution of blocks and their surrogates, this could even result in additional slowdowns.
To avoid such outcomes, we follow two rules when implementing speculation. First, we avoid enabling speculation on consecutive blocks to minimize accumulated errors and wasted resources. Second, speculation is not applied at the last block in the chain because the final output, once relayed back to BlockLLM’s scheduler, is not correctable. Thus, during serving, we selectively apply speculation to the top-$k$ bottleneck blocks—those characterized by their computation intensity or frequent use. The performance of speculative inference is evaluated in 7.3.
@@ 5.3 Request and Block Scheduling
BlockLLM’s request scheduling strategy prioritizes block instances that hold a request’s KV cache, provided the device memory can accommodate the request data. If the device memory cannot hold the data, the strategy estimates the latency of each candidate instance and schedules the request to the instance with the lowest latency increase. When a block does not have a request’s KV cache, the latency estimation for a candidate block instance on device $c$ is as follows:
$\begin{split}Latency_{d_{c}}&=T_{queue}+T_{compute}+T_{transfer}+T_{load},\\ T_{queue}&=\sum_{i=1}^{n}{Comp(req_{i})},\\ T_{compute}&=Comp(req),\\ T_{transfer}&=\begin{cases}T_{transfer}^{w/kv}&\text{agents dispatch,}\\ \dfrac{D_{req}}{B_{net}(s,d_{c})}&\text{scheduler dispatches,}\end{cases}\\ T_{load}&=\begin{cases}max(T_{swap}-T_{transfer},\ 0)\ &\text{idle device,}\\ T_{swap}&\text{loading overhead,}\\ \end{cases}\\ T_{swap}&=\dfrac{D_{b^{\prime}}}{B_{mem}(d_{c})}+\dfrac{D_{b}}{B_{net}(d_{c})}.\end{split}$
We incorporate four key factors: queuing ($T_{queue}$), computation ($T_{compute}$), transfer ($T_{transfer}$), and the overhead associated with block switching ($T_{load}$). $T_{queue}$ accounts for the duration required to process all $n$ queuing request batches. $T_{transfer}$ is subject to who initiates the dispatch: if done by the scheduler $s$, the request batch is at the start of the chain and no KV cache is involved. Otherwise, agent dispatching includes the costs of migrating the request and its KV cache (discussed in 5.1). Here, $D_{req}$ is the size of the request token, and $B_{net}(s,\ d_{j})$ denotes the network bandwidth between two devices. $T_{load}$ depends on the current status of the device. If the device is busy, this loading overhead only includes $T_{swap}$, the time to move out the existing block $b^{\prime}$ and load the new block $b$. If idle, the block’s loading can be overlapped with other operations. Therefore, the loading overhead is $T_{swap}-T_{transfer}$. In $T_{swap}$, $B_{mem}(d_{c})$ is the device memory bandwidth of candidate $d_{c}$; $D_{b^{\prime}}$ is the size of an existing block $b$. For simplicity, we keep block inside device memory and leave block swapping between other storage options for future work.
@@
In BlockLLM, the locality is quantified by monitoring historical traffic, specifically by a counter recording the frequency of being directly inter-dependent between two blocks. Block pairs with a higher locality value are placed within the same server. Furthermore, BlockLLM’s scheduler dynamically adapts to evolving traffic patterns; should the observed locality change, block instances are migrated as necessary to align with the new pattern. The benefit of locality-aware placement is discussed in 7.3.
6 Design Details
We have implemented a prototype of BlockLLM on top of vLLM. It is compatible with HuggingFace models. We use NCCL for data transfer among servers.
Profiling. To support the online serving system, BlockLLM profiles blocks by measuring computation time under different batch sizes, including surrogates and multiplexing performance. It also measures communication time between devices using NCCL primitives and the overhead of loading the block engine from disk to host and device memory.
Batching. Computation efficiency improves with larger batch sizes, but enforcing a fixed large batch size complicates request reorganization. Therefore, BlockLLM loosely encourages batching within each block instance. Upon receiving a new batch, BlockLLM’s agent queues it and attempts to pack it with neighboring requests, ensuring the combined batch does not exceed the upper batch size limit. If no requests are queued, the agent processes the batch directly. Requests reaching EOS are removed from the batch and sent to the scheduler.
Request dispatching. BlockLLM’s agents use a FIFO+priority queue, prioritizing requests that have left KV cache memory. Each block instance maintains a countdown clock for auto-regressive requests, prioritizing their return. The scheduler and agents dispatch requests differently: agents determine candidate blocks and pack them, broadcasting requests to available agents. The scheduler maintains a live record of block placements, simplifying dispatching.